Solve the following system of equations.
x + y = 3
2x + 2y = 6

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

The 2nd eqn is 2x the 1st.
Dependent
Infinite # of solutions.
No unique solution

Answer by stanbon(75887)

x + y = 3
2x + 2y = 6
------
Modify::
x + y = 3
x + y = 3
------
Solution: All points of the form (x,-x+3)
---------------
Cheers,
Stan H
===============